Solve the equation by expressing each side as a power of the same base and then simply exponents. Please help me solve this equation : 4^(3x+5)=1/256.
Answer by sudhanshu_kmr(1152)


2^2 =4 and 256 = 2^8

4^(3x+5)=1/256.
=> 2^(6x+10) = 2^(-8)
here value at base are equals, compare exponent..
6x+10 = -8
=> 6x = -18
=> x = -18/6 = -3
so, x = -3